to design a biasing current that is independent of power supply and temperature.

You will need to do both the circuit design and layout design of your project.

Design Targets:
The value of the output biasing current should be 25x10^-5 A.
You should ensure that the total current requirements of your circuit is as low as possible.

Operating Conditions
VCC: Min 2.7V, Typ 3.0V, Max 3.3V
Temperature: Min -45°C, Typ 25°C, Max 125°C
1
Expert's answer
2019-12-02T10:16:35-0500


In order to make a temperature-independent current source, you need to use a current mirror in which the circuit current is set by resistor 1, and resistor 2 is the load. It is necessary to find the resistance of resistor 1.Due to the fact that the base of the transistors are connected in both arms of the circuit, the same current flows.

The voltage across the resistor is:

"U_R=U_{dd}-0.7V=3V-0.7V=2.3V"

According to Ohm's law, we find the rotation R1:

"R_1=\frac{U_R}{I_R}=\frac{2.3V}{25\cdot 10^{-5}}=9.2 kOhm"